OpenStudy (anonymous):

A wise investor purchased a plot of land in 1947 for $84 000. In 1987 that same investor sold the land for $49 000 000? what annual rate of interest corresponds to an investment of $84 000 which grows to $49 000 000 in 40 years? (logarithms)

OpenStudy (anonymous):

so that is where it came from

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

not really a log problem

OpenStudy (anonymous):

i didn't think so but its art of the log q's in my book

OpenStudy (anonymous):

you want to solve \[4900=84(1+x)^{40}\] for x

OpenStudy (anonymous):

divide by 84 take the 40th root (with a calculator) subtract 1

OpenStudy (anonymous):

\[\frac{175}{3}=(1+x)^{40}\] \[\sqrt[40]{\frac{175}{3}}=1+x\] \[

OpenStudy (anonymous):

that dude made some crazy cash

OpenStudy (anonymous):

i get \[x+1=1.107\] \[x=.107=10.7\%\]

OpenStudy (anonymous):

thisis the answer in my book.. i just don't know how they got it.. \[49 000 000\div84 000000= (1+x)^{40}\] \[(1750 \div 3) ^{1/40) = 1+x (\[1730\div3^{1/40}\]-1=x 0.1726\[approxx\]

OpenStudy (anonymous):

i wrote it out yes? still confused let me know

OpenStudy (anonymous):

where do u get 175 and 3

OpenStudy (anonymous):

oh from reducing the fraction

OpenStudy (anonymous):

4900/84

OpenStudy (anonymous):

how do you do that

OpenStudy (anonymous):

i have forgotten all of my middle school math

OpenStudy (anonymous):

what i mean is how do u know when to stop reducing

OpenStudy (anonymous):

You can use logarithms to solve this A=P.e^rt A=value, P=original amount, r=interest rate, t=time, and e=e=2.71828 49million=84,000 x e^R.40 ln(49000/84)=ln(e^R.40) ln(49000/84)=40 x R R = 6.36/40 = 1.59 Look up "natural logarithm"

OpenStudy (anonymous):

When we didn't have calculators, finding the 40th roots of things was difficult. ;-)